A newspaper reported on the results of an opinion poll in which adults were asked what one thing they are most likely to do when they are home sick with a cold or the flu. In the survey, 61% said that they are most likely to sleep and 18% said that they would watch television. Although the sample size was not reported, typically opinion polls include approximately 1,000 randomly selected respondents. a. Assuming a sample size of 1,000 for this poll, construct a 95% confidence interval for the true percentage of all adults who would choose to sleep when they are at home The 95% confidence interval is (LIO). (Round to the nearest hundredth as needed.) b. If the true percentage of adults who would choose to sleep when they are at home sick is 71%, would you be surprised? O Yes O No
Answer #1

a)

Level of Significance,   α =    0.05
Sample Size,   n =    1000
      
Sample Proportion ,    p̂ =    0.61
z -value =Z0.05/2= 1.9600
Standard Error ,    SE = √[p̂(1-p̂)/n] =    0.0154
      
margin of error ,    E = Z*SE =    0.0302
      
Confidence Interval      
Interval Lower Limit , =    p̂ - E =    0.5798
Interval Upper Limit , =    p̂ + E =   0.6402

confidence interval is(0.58,0.64)

b)

yes,it would be a surprising result,as true percentage is outside the confidence interval